If k sin 2 x + (1/k) cosec 2 x = 2, for all x belonging to (0, pie/2) then cos 2 x + 5 sinx cosx + 6 sin 2 x is equals to (a) (k 2 +5k+6) / k 2 (b) (k 2 -5k+6) / k 2 (c)6 (d) none of these

If k sin2x + (1/k) cosec 2 x = 2, for all x belonging to (0, pie/2) then cos2x + 5 sinx cosx + 6 sin 2x is equals to 


 (a) (k2+5k+6) / k2  


(b) (k2-5k+6) / k2 


(c)6 


(d) none of these

Grade:12

1 Answers

Ramesh V
14 years ago

as k sin2x + (1/k) cosec 2 x = 2

which is (ksin2x - 1)2 = 0

or sin2x = 1/k  , cos2x = 1 -(1/k)

for cos2x + 5 sinx cosx + 6 sin 2x

= 5 - 5/k +6/k + 5/k*(k-1)1/2

= ( 1+5k+ 5*(k-1)1/2) / k

ans: none of the these

--

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and we will get you the answer and detailed solution very quickly.We are all IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Ramesh
IIT Kgp - 05 batch